OpenStudy (anonymous):

one number is 10 more than another number.Form a quadratic function that will allow you to find the minimum product of the two numbers. find the numbers and the minimum product.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

so far x=10+y how to make it into quadratic equation

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@calculusxy any idea

OpenStudy (anonymous):

Let \(x\) be the first number. Then the second number is \(x+10\). The product of the two is \(x(x+10)=x^2+10x\). To find the minimum of the product, you can complete the square to write this quadratic in vertex form \((x-a)^2+b\), which will be minimized when \(x=a\) since that makes the first term disappear, leaving you with \(b\).

OpenStudy (anonymous):

o wow

OpenStudy (anonymous):